Make New Entry,
Make Followup Entry
User name Markowitz
Log entry time 03:59:50 on July 08, 2010
Entry number 330758
keyword=Owl shift summary, part 1
Owl Shift Summary: July 7, 2010
- Bob Michaels
- Victor Maxwell
- Pete Markowitz
- 00:00 Accelerator is sending tune beam.
- 00:06 Move to BeO for beam location and spot size check. Run 1887
shows 8x6 raster is OK, and location is -6.5, 2.5.
- 00:22 Move to Ta for production running. Run 1848 is junk, with
all T1 triggers. Adjust prescales.
- 00:33 Start run 1849, production on Ta. Detector replay looks OK.
- 02:07 Start run 1850, production on Ta. Detector replay looks OK.
- 02:48 MCC calls to say a cavity has dropped out and they need to
re-LEM the machine.
- 02:53 Start run 1851, production on Ta.
- 03:04 MCC calls. "Zone 2L09-7 has dropped out of filaments." Beam
gone, end run 1851.
- 03:32 Start run 1852, production on Ta.